# Hybrid-z: Enhancing Kilo-Degree Survey bright galaxy sample photometric redshifts with deep learning

### v1

@@
Hybrid-z: Enhancing Kilo-Degree Survey bright galaxy sample
@@
We employ deep learning (DL) to improve photometric redshifts (photo-$z$s) in the Kilo-Degree Survey Data Release 4 Bright galaxy sample (KiDS-Bright DR4). This dataset, used as foreground for KiDS lensing and clustering studies, is flux-limited to $r<20$ mag with mean $z=0.23$ and covers 1000 deg2. Its photo-$z$s were previously derived with artificial neural networks from the ANNz2 package, trained on the Galaxy And Mass Assembly (GAMA) spectroscopy. Here we considerably improve over these previous redshift estimations by building a DL model, Hybrid-z, which combines an Inception-based convolutional neural network operating on four-band KiDS images with an artificial neural network using nine-band magnitudes from KiDS+VIKING. The Hybrid-z framework provides state-of-the-art photo-$z$s for KiDS-Bright, with negligible mean residuals of O($10^{-4}$) and scatter at the level of $0.014(1+z)$ – reduction by 20% over the previous nine-band derivations with ANNz2. Our photo-$z$s are robust and stable independently of galaxy magnitude, redshift, and color. In fact, for blue galaxies, which typically have more pronounced morphological features, Hybrid-z provides a larger improvement over ANNz2 than for red ones. We check our photo-$z$ model performance on test data drawn from GAMA, as well as from other KiDS-overlapping wide-angle spectroscopic surveys, namely SDSS, 2dFLenS and 2dFGRS. We find stable behavior and consistent improvement over ANNz2 throughout. We finally apply Hybrid-z trained on GAMA to the entire KiDS-Bright DR4 sample of 1.2 million galaxies. For these final predictions, we design a method of smoothing the input redshift distribution of the training set, to avoid propagation of features present in GAMA, related to its small sky area and large-scale structure imprint in its fields. Our work paves the way towards the best-possible photo-$z$s achievable with machine learning for any galaxy type both for the final KiDS-Bright DR5 data and for future deeper imaging, such as from the Legacy Survey of Space and Time.
@@ 1 Introduction
Redshift is a key quantity for cosmological analyses: as the basic proxy for galaxy distances, it allows us to map the large-scale structure of the Universe in time and three-dimensional space. Redshift can be measured to a sub-percent accuracy only via spectroscopy. For such spectroscopic redshifts (spec-$z$s) we first collect the spectrum of an object and then identify the shift in spectral lines with respect to the rest frame. However, even in the current era of fast measurements, such as with the Dark Energy Spectroscopic Instrument (DESI Collaboration et al., 2016), we are able to obtain spec-$z$s only for a small fraction of all detected galaxies.
@@
On the other hand, we can estimate redshifts for a much larger sample from photometric measurements. Such photometry-based redshifts (photo-$z$) are an alternative method based on the correlation between redshift and galaxy apparent magnitudes, originally pointed out by Baum (1957) and first applied for photoelectric magnitudes in nine pass-bands by Baum (1962). Two main attractive factors of photo-$z$s are that they allow us to obtain redshifts for galaxies fainter than generally possible with spectroscopy; and that the number of objects with redshift estimates per unit telescope time is also much larger (Hildebrandt et al., 2010). However, photo-$z$s cannot provide as good redshift accuracy and precision as spec-$z$s; they are, nevertheless, indispensable in today’s massive imaging surveys cataloging millions and billions of galaxies.
@@
Photo-$z$s are based on a complicated mapping from photometry to redshift space, and they are difficult to handle analytically since this mapping depends on observational, computational, and statistical factors. Photo-z estimation methods can be generally categorized into template fitting and empirical approaches. In the latter case, which is our focus in this paper, the relation between photometric quantities and redshift is very commonly found by machine learning (ML) algorithms, although we note that simpler approaches of using functional fitting have also been proposed (e.g. Connolly et al., 1995; Krone-Martins et al., 2014).
@@
In supervised ML techniques, the algorithm derives the empirical relation between observed quantities and labels from appropriate training on labeled data. Therefore, their main challenge and limitation lies in extrapolating the results beyond a representative training set. However, if such appropriate training data exist, ML methods can excel and this led to proposing numerous approaches for photo-$z$s based on the newest developments in computer science. Some examples, among others, include support vector machines (Wadadekar, 2004), random forest (Carliles et al., 2010; Li et al., 2021), artificial neural networks (ANN; e.g. Tagliaferri et al., 2003; Collister & Lahav, 2004; Oyaizu et al., 2008), ensemble learning (Cunha & Humphrey, 2022), Gaussian processes (Way & Srivastava, 2006; Bonfield et al., 2010), self-organizing maps (Way & Klose, 2012), k-nearest neighbors (Graham et al., 2017), mixture density networks (Ansari et al., 2021), and finally deep neural networks (e.g. Hoyle, 2016; D’Isanto & Polsterer, 2018).
@@
Among the various supervised ML techniques for photo-$z$ derivation, deep learning (DL) has emerged as a particularly promising one. DL makes it possible to entirely skip ‘higher-level’ quantities such as galaxy magnitudes or sizes derived from photometric post-processing, and build the ML model using multi-band imaging directly. In such frameworks, ‘deep’ indicates that the models usually have compounded multi-layer architectures. Their usage for photo-$z$s was pioneered by Hoyle (2016) and then studied by e.g. D’Isanto & Polsterer (2018); Menou (2019); Pasquet et al. (2019); Dey et al. (2022); Henghes et al. (2022); Treyer et al. (2024) for Sloan Digital Sky Survey (SDSS) data, Schuldt et al. (2021) for Hyper-Suprime Cam (HSC), by Li et al. (2022) for the Kilo-Degree Survey (KiDS) or by Roster et al. (2024) for DESI Imaging. These papers have demonstrated that using images directly, or in combination with magnitudes, allows one to derive photo-$z$s of better performance than those based on summary galaxy data such as magnitudes.
@@
A particular realization of DL is convolutional neural networks (CNN). They are a type of ANN suitable for computer vision problems, such as image feature detection or classification. CNN is inspired by the human vision system. Similar to real neurons, which receive input and pass electrochemical signals (McCulloch, 1943), artificial neurons are also used in CNNs. The significance of the CNN is evidenced by their vast use in image recognition tasks (e.g. LeCun et al., 1998). CNN is appropriate for processing data that have grid-like topology, such as images, because of its local connectivity, parameter sharing, and translational invariance. We chose CNNs to extract the galaxy image patterns by detecting features like edges, textures, and shapes, which is expected to improve photo-$z$ derivations over the methods not employing such information.
@@
In this work, we extend the previous feature-based ML efforts to build a successful DL model, which we call ‘Hybrid-z’, that integrates both images and features for KiDS-Bright photo-$z$ estimations. We use the same GAMA training set as in B21 and construct a photo-z model that is conceptually similar to one of those tested in Li et al. (2022). Namely, it combines a deep convolutional network, employing $ugri$ imaging, with an ordinary ANN that is fed by 9-band galaxy magnitudes. An analogous configuration was also studied by Henghes et al. (2022) for SDSS, and inspired by their results, we also use ‘Inception’ as our basic architecture for the DL part.
This paper is organized in the following manner. In Sect. 2 we describe the data used, next, in Sect.  3 we explain the basic concepts of CNNs and a special CNN architecture that we used, called Inception. In Sect. 4, we describe our Hybrid-z model to estimate the photo-$z$s and the statistics that we use to measure the performance of the network. Sect. 5 presents our results and in Sect. 6, we conclude and discuss future prospects.
@@ 2 Data
In this section, we discuss the data used in this study. We have employed the KiDS DR4 Bright sample images supplemented with photometry (i.e., magnitudes). The required training and testing data are labeled using the spectroscopic redshifts from the GAMA survey.
@@ 2.1 KiDS images and photometry
KiDS is an optical wide-field imaging survey of the European Southern Observatory (ESO) at the VLT Survey Telescope (VST, Capaccioli & Schipani, 2011), having at the focal plane a 268 million pixel CCD mosaic camera called OmegaCAM (Kuijken, 2008). VST is an alt-az mounted modified Ritchey-Cretien telescope located in the ESO Paranal Observatory, Chile. The images were taken in four broad bands ($ugri$), and the survey covers 1350 square degrees of the extragalactic sky. The final footprint of the survey is shown in fig. 3 of Wright et al. (2024); here we use its publicly available subset.
@@
KiDS-ESO Data Release 4, shortly KiDS DR4 (Kuijken et al., 2019), is the fourth public release of KiDS. The ASTRO-WISE optical pipeline and data reduction environment (McFarland et al., 2013) is used to produce stacked (or co-added) composite images created by combining multiple individual exposures of the same sky area, one in each of the four bands. As a result, KiDS DR4 optical data are organized into $4\times 1006$ one square-degree tiles.
@@
KiDS DR4 products consist of astrometrically and photometrically calibrated co-added images with a uniform pixel scale of 0.2 arcsec. The pixel units are fluxes relative to the 0th magnitude (de Jong et al., 2015). We have downloaded the KiDS DR4 tiles [^1] https://kids.strw.leidenuniv.nl/DR4 and made cutouts of galaxies with a size of $7.2^{\prime\prime}\times 7.2^{\prime\prime}$ ($36\times 36$ pixels), since most of the objects we use are smaller than this. We have also tried bigger cutouts like $20.2^{\prime\prime}\times 20.2^{\prime\prime}$ and $8^{\prime\prime}\times 8^{\prime\prime}$, but the smaller size reduces the noise in images without losing galaxy flux information. Too large cutouts could also lead to frequent situations when more than one galaxy appears in the image. This type of contamination could adversely affect the performance of the model, although it has been argued in the literature that CNNs for photo-z estimation could in fact benefit from physically close pairs in the images (e.g. Pasquet et al., 2019). In our case we also use magnitudes of individual galaxies, which should mitigate this effect, be it positive or negative. In any case, for our fiducial cutout size, more than one object is present in the image very rarely, in less than 1% cases.
@@
Finally, we normalized the pixel values to the range $[0,1]$ galaxy-wise, i.e. jointly for all the $ugri$ bands for a given galaxy cutout. The normalization formula is:
$X_{normalized}=\frac{X-X_{min}}{X_{max}-X_{min}}$ (1)
@@
The minimum value of the pixels for the 4-band per-galaxy images (cutouts) is denoted as $X_{min}$ and the maximum value as $X_{max}$. In occasional cases when the image has (a) saturated pixel(s), this normalization will not work properly. Dealing with this problem would be beyond the scope of our work as it would require building an extra framework to pre-analyze all the 4.8 million cutouts and cleaning them up of such artifacts. We note that in most cases, objects located in such corrupted areas will not be useful for science anyway, as they will bear an appropriate KiDS MASK value indicating that their photometry is not reliable (see next Section). According to our estimates, for the ‘clean’ data (i.e. those not affected by the mask), the presence of artifacts in the images is very infrequent, at a fraction of a percent level. Last but not least, the usage of magnitudes together with images in the model, will minimize their influence on the derived photo-zs.
@@
We standardized these 9-band magnitude features using the StandardScaler class from the scikit-learn Python library (Pedregosa et al., 2011), which computes the mean ($\overline{m}$) and standard deviation ($\sigma_{m}$) for each band independently. The magnitude values ($m$) in each band are transformed as:
$m_{standardized}=\frac{m-\overline{m}}{\sigma_{m}}$ (2)
@@
2.2 KiDS-Bright Galaxy Sample
In this work, we derive DL photo-$z$s for the KiDS DR4 Bright Galaxy Sample, introduced in B21. This dataset contains galaxies selected from KiDS DR4 with the flux limit $r_{\mathrm{auto}}<20$ mag, where ‘auto’ stands for SExtractor-derived estimate of the total flux via automatic aperture photometry (Bertin & Arnouts, 1996). The selection of the sample is designed to have the best possible match with overlapping GAMA-equatorial spectroscopic data, which originally were selected using SDSS Petrosian magnitudes in the $r$ band (Liske et al., 2015). As it was discussed in B21, among the various $r$-band magnitude measurements in KiDS, a cut in $r_{\mathrm{auto}}$ provides the best correspondence to the GAMA original $r_{\mathrm{Petro}}<19.8$ limit.

### v2

# Hybrid-z: Enhancing the Kilo-Degree Survey bright galaxy sample photometric redshifts with deep learning

@@
Hybrid-z: Enhancing the Kilo-Degree Survey bright galaxy sample
@@
We employed deep learning to improve the photometric redshifts (photo-$z$s) in the Kilo-Degree Survey Data Release 4 bright galaxy sample (KiDS-DR4 Bright). This dataset, used as foreground for KiDS lensing and clustering studies, is flux-limited to $r<20$ mag with mean $z=0.23$ and covers 1000 deg2. Its photo-$z$s were previously derived with artificial neural networks from the ANNz2 package trained on the Galaxy And Mass Assembly (GAMA) spectroscopy. Here, we considerably improve on these previous redshift estimations by building a deep learning model, Hybrid-z, that combines an inception-based convolutional neural network operating on four-band KiDS images with an artificial neural network using nine-band magnitudes from KiDS+VIKING. The Hybrid-z framework provides state-of-the-art photo-$z$s for KiDS-Bright with negligible mean residuals of O($10^{-4}$) and scatter at a level of $0.014(1+z)$ – representing a reduction of 20% compared to the previous nine-band derivations with ANNz2. Our photo-$z$s are robust and stable independently of galaxy magnitude, redshift, and color. In fact, for blue galaxies, which typically have more pronounced morphological features, Hybrid-z provides a larger improvement over ANNz2 than for red galaxies. We checked our photo-$z$ model performance on test data drawn from GAMA as well as from other KiDS-overlapping wide-angle spectroscopic surveys, namely SDSS, 2dFLenS, and 2dFGRS. We found stable behavior and consistent improvement over ANNz2 throughout. Finally, we applied Hybrid-z trained on GAMA to the entire KiDS-Bright DR4 sample of 1.2 million galaxies. For these final predictions, we designed a method of smoothing the input redshift distribution of the training set in order to avoid propagation of features present in GAMA related to its small sky area and large-scale structure imprint in its fields. Our work paves the way toward the best-possible photo-$z$s achievable with machine learning for any galaxy type for both the final KiDS-Bright DR5 data and for future deeper imaging, such as from the Legacy Survey of Space and Time.
@@ 1 Introduction
Redshift is a key quantity for cosmological analyses. As the basic proxy for galaxy distances, it allows one to map the large-scale structure of the Universe in time and three-dimensional space. Redshift can be measured to a sub-percent accuracy only via spectroscopy. For such spectroscopic redshifts (spec-$z$s), one first collects the spectrum of an object and then identifies the shift in spectral lines with respect to the rest frame. However, even in the current era of fast measurements, such as with the Dark Energy Spectroscopic Instrument (DESI Collaboration et al., 2016), spec-$z$s can be obtained only for a small fraction of all detected galaxies.
@@
On the other hand, redshifts can be estimated for a much larger sample from photometric measurements. Such photometry-based redshifts (photo-$z$) provide an alternative method to the spec-$z$s and are based on the correlation between redshift and apparent galaxy magnitudes. This approach was originally pointed out by Baum (1957) and first applied to obtain photoelectric magnitudes in nine passbands by Baum (1962). The two main attractive factors of photo-$z$s are that they allow one to obtain redshifts for galaxies fainter than generally possible with spectroscopy and that the number of objects with redshift estimates per unit telescope time is also much larger (Hildebrandt et al., 2010). Notably, photo-$z$s cannot provide redshift accuracy and precision as good as spec-$z$s. Nevertheless, they are indispensable in today’s massive imaging surveys cataloging millions and billions of galaxies.
@@
Photo-$z$s are based on a complicated mapping from photometry to redshift space, and they are difficult to handle analytically since this mapping depends on observational, computational, and statistical factors. Photo-z estimation methods can be generally categorized into template fitting and empirical approaches. In the latter case, which is our focus in this paper, the relation between photometric quantities and redshift is very commonly found by machine learning (ML) algorithms, although we note that simpler approaches of using functional fitting have also been proposed (e.g., Connolly et al., 1995; Krone-Martins et al., 2014).
@@
In supervised ML techniques, the algorithm derives the empirical relation between observed quantities and labels from appropriate training on labeled data. Therefore, their main challenge and limitation lies in extrapolating the results beyond a representative training set. However, if such appropriate training data exist, ML methods can excel, and this has led to numerous approaches being proposed for photo-$z$s based on the newest developments in computer science. Some examples include support vector machines (Wadadekar, 2004), random forest (Carliles et al., 2010; Li et al., 2021), artificial neural networks (ANNs; e.g., Tagliaferri et al., 2003; Collister & Lahav, 2004; Oyaizu et al., 2008), ensemble learning (Cunha & Humphrey, 2022), Gaussian processes (Way & Srivastava, 2006; Bonfield et al., 2010), self-organizing maps (Way & Klose, 2012), k-nearest neighbors (Graham et al., 2017), mixture density networks (Ansari et al., 2021), and finally deep neural networks (e.g., Hoyle, 2016; D’Isanto & Polsterer, 2018).
@@
Among the various supervised ML techniques for photo-$z$ derivation, deep learning (DL) has emerged as a particularly promising one. Deep learning makes it possible to entirely skip ”higher-level” quantities such as galaxy magnitudes or sizes derived from photometric post-processing and build the ML model using multi-band imaging directly. In such frameworks, ”deep” indicates that the models usually have compounded multi-layer architectures. Their usage for photo-$z$s was pioneered by Hoyle (2016) and then studied by, for example, D’Isanto & Polsterer (2018); Menou (2019); Pasquet et al. (2019); Dey et al. (2022); Henghes et al. (2022); Treyer et al. (2024) for Sloan Digital Sky Survey (SDSS) data; Schuldt et al. (2021) for Hyper-Suprime Cam (HSC); Li et al. (2022) for the Kilo-Degree Survey (KiDS); and by Roster et al. (2024) for DESI Imaging. These papers have demonstrated that using images directly, or in combination with magnitudes (e.g., Li et al., 2022; Jones et al., 2024; Roster et al., 2024), allows one to derive photo-$z$s of better performance than those based on tabular galaxy data such as magnitudes.
@@
A particular realization of DL is convolutional neural networks (CNNs). They are a type of ANN suitable for computer vision problems, such as image feature detection or classification, and are inspired by the human vision system. Similar to real neurons, which receive input and pass electrochemical signals (McCulloch, 1943), artificial neurons are used in CNNs. The significance of CNNs is evidenced by their vast use in image recognition tasks (e.g., LeCun et al., 1998). These networks are appropriate for processing data that have grid-like topology, such as images, because of their local connectivity, parameter sharing, and translational invariance. We chose CNNs to extract the galaxy image patterns by detecting features such as edges, textures, and shapes, which are expected to improve photo-$z$ derivations compared to methods that do not employ such information.
In this paper, we present a photo-z estimation in which we incorporate both fluxes (magnitudes) and multichannel galaxy images by using DL techniques for KiDS (de Jong et al., 2013). KiDS is a multiband imaging survey covering about 1350 deg2 of the sky, of which we employ $\sim 1000$ deg2 from its fourth data release (Kuijken et al., 2019). The DL photo-$z$s within KiDS have already been studied in detail by Li et al. (2022), where various setups using both images and magnitudes were compared for a general selection of galaxies spanning $0<z\lesssim 3$. However, until now, such imaging-based photo-$z$ approaches have not been employed in KiDS in the regime where they are expected to bring the most improvement over ”shallow” ML, namely, for relatively bright and well-resolved galaxies. Here, we fill this gap and focus only on the bright end of the KiDS data.
@@
In this work, we extend the previous feature-based ML efforts to build a successful DL model that we call ”Hybrid-z.” The model integrates both images and features for KiDS-DR4 Bright photo-$z$ estimations. We used the same GAMA training set as in B21 and constructed a photo-z model that is conceptually similar to one tested in Li et al. (2022). Namely, it combines a deep convolutional network, employing $ugri$ imaging, with an ordinary ANN that is fed by nine-band galaxy magnitudes. An analogous configuration was also studied by Henghes et al. (2022) for SDSS, and inspired by their results, we also use ”inception” as our basic architecture for the DL part.
This paper is organized in the following manner. In Sect. 2 we describe the data used. Next, in Sect.  3 we explain the basic concepts of CNNs and the special CNN architecture that we used, called inception. In Sect. 4, we describe our Hybrid-z model to estimate the photo-$z$s and the statistics that we used to measure the performance of the network. Sect. 5 presents our results, and in Sect. 6, we conclude and discuss future prospects.
@@ 2 Data
In this section, we discuss the data used in this study. We employed the KiDS-Bright DR4 sample images supplemented with photometry (i.e., magnitudes). The required training and testing data are labeled using the spectroscopic redshifts from the GAMA survey.
@@ 2.1 KiDS images and photometry
Kilo-degree survey is an optical wide-field imaging survey of the European Southern Observatory (ESO) at the Very Large Telescope (VLT) Survey (VST, Capaccioli & Schipani, 2011), having at the focal plane a 268 million pixel Charge-Coupled Device (CCD) mosaic camera called OmegaCAM (Kuijken, 2008). VST is an alt-az mounted modified Ritchey-Cretien telescope located in the ESO Paranal Observatory, Chile. The images were taken in four broad bands ($ugri$), and the survey covers 1350 square degrees of the extragalactic sky. The final footprint of the survey is shown in fig. 3 of Wright et al. (2024); here we use its publicly available subset.
@@
KiDS-ESO Data Release 4 (KiDS DR4; (Kuijken et al., 2019)), is the fourth public release of KiDS. The ASTRO-WISE optical pipeline and data reduction environment (McFarland et al., 2013) is used to produce stacked (or co-added) composite images created by combining multiple individual exposures of the same sky area, one in each of the four bands. As a result, KiDS DR4 optical data are organized into $4\times 1006$ one square-degree tiles.
KiDS DR4 products consist of astrometrically and photometrically calibrated co-added images with a uniform pixel scale of 0.2 arcsec. The pixel units are fluxes relative to the 0th magnitude (de Jong et al., 2015). We have downloaded the KiDS DR4 tiles [^1] https://kids.strw.leidenuniv.nl/DR4 and made cutouts of galaxies with a size of $7.2^{\prime\prime}\times 7.2^{\prime\prime}$ ($36\times 36$ pixels), since most of the objects we use are smaller than this. In particular, this cutout size is above the 99-percentile level of the half-light diameter (i.e., 2 $\times$ FLUX_RADIUS) in KiDS-Bright. The largest galaxies, not fitting within our cutouts, are likely of little interest for our work anyway: they will be very nearby and will have had spec-$z$ from wide-angle surveys or otherwise will not be useful for lensing studies.
@@
We have also tried bigger cutouts such as $20.2^{\prime\prime}\times 20.2^{\prime\prime}$ and $8^{\prime\prime}\times 8^{\prime\prime}$, motivated receptively by Grespan et al. (2024) and Li et al. (2022), but our finally adopted size gave the best results. The smaller size reduces the noise in images without losing galaxy flux information. Too large cutouts could also lead to frequent situations when more than one galaxy appears in the image. This type of contamination could adversely affect the performance of the model, although it has been argued in the literature that CNNs for photo-z estimation could in fact benefit from physically close pairs in the images (e.g., Pasquet et al., 2019). In our case we also use magnitudes of individual galaxies, which should mitigate this effect, be it positive or negative. In any case, for our fiducial cutout size, more than one object is present in the image very rarely, in less than 1% cases.
@@
Finally, we normalized the pixel values to the range $[0,1]$ galaxy-wise, i.e. jointly for all the $ugri$ bands for a given galaxy cutout. The normalization formula is
$X_{normalized}=\frac{X-X_{min}}{X_{max}-X_{min}}.$ (1)
@@
The minimum value of the pixels for the 4-band per-galaxy images (cutouts) is denoted as $X_{min}$ and the maximum value as $X_{max}$. In occasional cases when the image has (a) saturated pixel(s), this normalization will not work properly. Dealing with this problem would be beyond the scope of our work as it would require building an extra framework to pre-analyze all the 4.8 million cutouts and cleaning them up of such artifacts. We note that in most cases, objects located in such corrupted areas will not be useful for science anyway, as they will bear an appropriate KiDS MASK value indicating that their photometry is not reliable (see next Section). According to our estimates, for the ‘clean’ data (i.e., those not affected by the mask), the presence of artifacts in the images is very infrequent, at a fraction of a percent level. Last but not least, the usage of magnitudes together with images in the model, will minimize their influence on the derived photo-zs.
@@
We standardized these 9-band magnitude features using the StandardScaler class from the scikit-learn Python library (Pedregosa et al., 2011), which computes the mean ($\overline{m}$) and standard deviation ($\sigma_{m}$) for each band independently. The magnitude values ($m$) in each band are transformed as
$m_{standardized}=\frac{m-\overline{m}}{\sigma_{m}}.$ (2)
@@
2.2 KiDS-DR4 Bright galaxy sample
In this work, we derive DL photo-$z$s for the KiDS-DR4 Bright galaxy sample, introduced in B21. This dataset contains galaxies selected from KiDS DR4 with the flux limit $r_{\mathrm{auto}}<20$ mag, where ”auto” stands for SExtractor-derived estimate of the total flux via automatic aperture photometry (Bertin & Arnouts, 1996). The selection of the sample is designed to have the best possible match with overlapping GAMA-equatorial spectroscopic data, which originally were selected using SDSS Petrosian magnitudes in the $r$ band (Liske et al., 2015). As it was discussed in B21, among the various $r$-band magnitude measurements in KiDS, a cut in $r_{\mathrm{auto}}$ provides the best correspondence to the GAMA original $r_{\mathrm{Petro}}<19.8$ limit.